# Characterizing the minimax rate of nonparametric regression under bounded convex constraints

### v3

@@ 4       $k\leftarrow k+1$;
Now observe that $\bigcup_{J\leq J^{\ast}}[\varepsilon_{J},\varepsilon_{J-1})=[\varepsilon_{J^{\ast}},\infty)$. Pick any $x\geq\varepsilon_{J^{\ast}}$. Then $x\in[\varepsilon_{J},\varepsilon_{J-1})$ for some $J\leq J^{\ast}$. Thus
$2\kappa x\geq 2\kappa\cdot\varepsilon_{J}=\kappa\cdot\varepsilon_{J-1}.$
Note further that $x\mapsto\exp(-C^{\prime}nx^{2})$ is non-increasing for $x\geq 0$. Consequently,
$\mathbb{P}_{\bar{f}}\left(\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}\geq 2\kappa x\right)$ $\leq\mathbb{P}_{\bar{f}}\left(\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}\geq\kappa\cdot\varepsilon_{J-1}\right)$
$\leq\underline{C}\cdot\mathbbm{1}(J^{\ast}>1)\exp(-C^{\prime}n\varepsilon_{J-1}^{2})$
$\leq\underline{C}\cdot\mathbbm{1}(J^{\ast}>1)\exp(-C^{\prime}nx^{2}).$
This proves that $\mathbb{P}_{\bar{f}}\left(\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}\geq 2\kappa x\right)\leq\underline{C}\cdot\mathbbm{1}(J^{\ast}>1)\exp(-C^{\prime}nx^{2})$ for all $x\geq\varepsilon_{J^{\ast}}$.
Then we compute $\mathbb{E}_{X,Y}\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}^{2}$ by integration, splitting the integral into two parts:
$\mathbb{E}_{X,Y}\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}^{2}$ $=\int_{0}^{\infty}2x\mathbb{P}_{\bar{f}}\left(\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}\geq x\right)\mathrm{d}x$
$\leq\int_{0}^{2\kappa\varepsilon_{J^{\ast}}}2x\mathrm{d}x+\int_{2\kappa\varepsilon_{J^{\ast}}}^{\infty}2x\mathbb{P}_{\bar{f}}\left(\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}\geq x\right)\mathrm{d}x$
$\leq C^{\prime\prime\prime}{\varepsilon_{J^{\ast}}}^{2}+\int_{2\kappa\varepsilon_{J^{\ast}}}^{\infty}2x\cdot\underline{C}\cdot\mathbbm{1}(J^{\ast}>1)\exp(-C^{\prime\prime}nx^{2})\mathrm{d}x$
$=C^{\prime\prime\prime}{\varepsilon_{J^{\ast}}}^{2}+C^{\prime\prime\prime\prime}\cdot\mathbbm{1}(J^{\ast}>1)n^{-1}\exp(-C^{\prime\prime\prime\prime\prime}n{\varepsilon_{J^{\ast}}}^{2}).$
Recall from our assumption (2.6) that either $J^{\ast}=1$ or $n{\varepsilon_{J^{\ast}}}^{2}>\log 2$. In either case, $\mathbb{E}_{X,Y}\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}^{2}\leq\overline{C}{\varepsilon_{J^{\ast}}}^{2}$ for some constant $\overline{C}$ depending only on $\sigma$, $B_{\mathcal{F}}$, and our constant $C>3$.
Theorem 2.12.
Define $\varepsilon^{\ast}:=\sup\{\varepsilon:n\varepsilon^{2}\leq\log M_{\mathcal{F}}^{\operatorname{loc}}(\varepsilon,c)\}$ where $c$ is a sufficiently large absolute constant. Then the minimax rate is ${\varepsilon^{\ast}}^{2}\wedge d^{2}$ up to absolute constants.
{toappendix}
@@
Suppose $n{\varepsilon^{\ast}}^{2}>8\log 2$. Set $\alpha=\min(\sigma/2,1)\in(0,1]$ and $\delta^{\ast}=\alpha\varepsilon^{\ast}$. Note that $\alpha^{2}\leq\frac{\sigma^{2}}{4}$. Then using the monotonicity of $M_{\mathcal{F}}^{\operatorname{loc}}(\cdot,c)$, we obtain
$\log M_{\mathcal{F}}^{\operatorname{loc}}(\delta^{\ast},c)$ $\geq\log M_{\mathcal{F}}^{\operatorname{loc}}(\varepsilon^{\ast},c)\geq n{\varepsilon^{\ast}}^{2}=\frac{n{\varepsilon^{\ast}}^{2}}{2}+\frac{n{\varepsilon^{\ast}}^{2}}{2}>\frac{n{\delta^{\ast}}^{2}}{2\alpha^{2}}+4\log 2$
$\geq 4\cdot\frac{n{\delta^{\ast}}^{2}}{2\sigma^{2}}+4\log 2\geq 4\left(\frac{n{\delta^{\ast}}^{2}}{2\sigma^{2}}\vee\log 2\right).$
@@
Thus, the lower bound condition from Lemma 2.3 is satisfied for $\delta^{\ast}$, so indeed ${\varepsilon^{\ast}}^{2}$ is a lower bound for the minimax rate up to absolute constants by the lemma when $n{\varepsilon^{\ast}}^{2}>8\log 2$.
@@
We now show the minimax rate has an upper bound of rate ${\varepsilon^{\ast}}^{2}$ when $n{\varepsilon^{\ast}}^{2}>8\log 2$. Following lecamshamindramatey2022, we set $\eta=\frac{c\sqrt{2}}{L(c/2-1,\sigma,B_{\mathcal{F}})}\wedge 1\in(0,1)$, pick $D$ sufficiently large such that $D\eta>1$, and set $\delta=D\sqrt{2}\varepsilon^{\ast}$. Again using monotonicity of $M_{\mathcal{F}}^{\operatorname{loc}}(\cdot,c)$ and the definition of $\varepsilon^{\ast}$, we have
$n\delta^{2}$ $=2D^{2}n{\varepsilon^{\ast}}^{2}=\frac{2}{\eta^{2}}\cdot n(D\eta\cdot\varepsilon^{\ast})^{2}>\frac{2}{\eta^{2}}\cdot\log M_{\mathcal{F}}^{\operatorname{loc}}(D\eta\varepsilon^{\ast},c)\geq 2\log M_{\mathcal{F}}^{\operatorname{loc}}(D\eta\varepsilon^{\ast},c)$
@@
$\geq 2\log M_{\mathcal{F}}^{\operatorname{loc}}\left(\frac{c\sqrt{2}\cdot D\varepsilon^{\ast}}{L(c/2-1,\sigma,B_{\mathcal{F}})},c\right)=2\log M_{\mathcal{F}}^{\operatorname{loc}}\left(\frac{c\cdot\delta}{L(c/2-1,\sigma,B_{\mathcal{F}})},c\right)$
Further, since $n{\varepsilon^{\ast}}^{2}>8\log 2$, we have $n\delta^{2}=2nD^{2}{\varepsilon^{\ast}}^{2}>16D^{2}\log 2>\log 2$, so indeed
@@
$n\delta^{2}>2\log M_{\mathcal{F}}^{\operatorname{loc}}\left(\frac{\delta\cdot c}{\sqrt{L(c/2-1,\sigma,B_{\mathcal{F}})}},c\right)\vee\log 2.$
In other words, $\delta$ satisfies the condition in (2.6).
Now consider the map
@@
$\psi\colon 0<x\mapsto nx^{2}-2\log M_{\mathcal{F}}^{\operatorname{loc}}\left(x\cdot\frac{c}{\sqrt{L(c/2-1,\sigma,B_{\mathcal{F}})}},c\right)\vee\log 2$
@@
which is non-decreasing. We have shown that $\psi(\delta)>0$. Recall our definition of $\varepsilon_{J^{\ast}}$ from Theorem 2.11, and let us show that $\delta\geq\varepsilon_{J^{\ast}}/2$. Suppose no $J\geq 1$ exists such that $\varepsilon_{J}$ satisfies (2.6), in which case $J^{\ast}=1$. Then $\psi(\varepsilon_{J^{\ast}}/2)\leq\psi(\varepsilon_{J^{\ast}})<0<\psi(\delta)$ so $\delta\geq\varepsilon_{J^{\ast}}/2$. On the other hand, suppose some $J$ does exist such that $\varepsilon_{J}$ satisfies (2.6). Because $J^{\ast}<\infty$ is chosen maximally in our algorithm, $\varepsilon_{J^{\ast}+1}=\varepsilon_{J^{\ast}}/2$ does not satisfy (2.6), so we have $\psi(\varepsilon_{J^{\ast}}/2)=\psi(\varepsilon_{J^{\ast}+1})<0<\psi(\delta)$, so $\delta\geq\varepsilon_{J^{\ast}}/2$.
@@
Since the minimax rate has an upper bound of $\varepsilon_{J^{\ast}}^{2}$ up to constants by Theorem 2.11, it follows that the rate is also upper bounded by $4\delta^{2}=8D^{2}{\varepsilon^{\ast}}^{2}$. Thus, the minimax has a lower and upper bound of rate ${\varepsilon^{\ast}}^{2}$ up to absolute constants. But if ${\varepsilon^{\ast}}^{2}$ is a lower bound, then so is ${\varepsilon^{\ast}}^{2}\wedge d$; further, the squared diameter is always an upper bound for the minimax rate, so ${\varepsilon^{\ast}}^{2}\wedge d^{2}$ is also an upper bound up to absolute constants. Thus, if $n{\varepsilon^{\ast}}^{2}>8\log 2$, the minimax rate is ${\varepsilon^{\ast}}^{2}\wedge d^{2}$ up to absolute constants.
Next, we consider the case where $n{\varepsilon^{\ast}}^{2}\leq 8\log 2$. Observe then that
$\log M_{\mathcal{F}}^{\operatorname{loc}}(2\varepsilon^{\ast},c)\leq 4n{\varepsilon^{\ast}}^{2}\leq 32\log 2.$
@@
Now take two endpoints of a diameter of $\mathcal{F}$ and form a ball $B$ of radius $2\varepsilon^{\ast}$ centered at the midpoint of these two endpoints. Consider the diameter of $B$ that coincides with the chosen diameter of $\mathcal{F}$ at infinitely many points. Since an interval of length $4\varepsilon^{\ast}$ if partitioned into subintervals of length less than $(2\varepsilon^{\ast}/c)$ is composed of at least $2c$ subintervals, for sufficiently large $c$, at least $\exp(32\log 2)$ many points at least $(2\varepsilon^{\ast}/c)$ distance apart can be chosen along this particular diameter of $B$. All of these points cannot also belong to $\mathcal{F}$, otherwise we will have formed a $(2\varepsilon^{\ast}/c)$-packing of $B\cap\mathcal{F}$ with $\exp(32\log 2)$ points, contradicting $\log M_{\mathcal{F}}^{\operatorname{loc}}(2\varepsilon^{\ast},c)\leq 32\log 2$. This means the endpoints of the diameter of $\mathcal{F}$ are strictly contained in the diameter of $B$ we picked, so that $d\leq 4\varepsilon^{\ast}\leq 8\sqrt{2n^{-1}\log 2}$.
@@
Now, consider an $\varepsilon$ which is proportional to $d$. In this case, $n\varepsilon^{2}$ is upper bounded by some constant since $d\leq 8\sqrt{2n^{-1}\log 2}$, while $\log M_{\mathcal{F}}^{\operatorname{loc}}(2\varepsilon,c)$ can be made arbitrarily large by increasing $c$, to ensure the condition from Lemma 2.3 for the lower bound is attained. So $\varepsilon^{2}$ is a lower bound for the minimax-rate up to constants, and thus by proportionality, $d^{2}$ is a lower bound up to absolute constants. Thus ${\varepsilon^{\ast}}^{2}\wedge d^{2}$ is a lower bound for the minimax rate.

### v5

@@ Appendix B Proofs for Section 3
Pick any $x\geq\varepsilon_{J^{\ast}}$. Then for some integer $-\infty<J\leq J^{\ast}$, $x\in[\varepsilon_{J},\varepsilon_{J-1})$ since $\bigcup_{J\leq J^{\ast}}[\varepsilon_{J},\varepsilon_{J-1})=[\varepsilon_{J^{\ast}},\infty)$. We have by Lemma 3.3 and definitions of $\kappa$, $\varepsilon_{J}$, $f^{\ast}$, and $f^{\ast\ast}$ that
$\|\bar{f}-f^{\ast\ast}\|_{L_{2}(\mathbb{P}_{X})}$ $\leq\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}+\|f^{\ast}-f^{\ast\ast}\|_{L_{2}(\mathbb{P}_{X})}$
$=\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}+\|\Upsilon_{J^{\ast}}-\Upsilon_{J^{\ast\ast}+1}\|_{L_{2}(\mathbb{P}_{X})}$
$\leq\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}+\tfrac{d(2+4c)}{c2^{J^{\ast}}}$
$=\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}+\tfrac{1+2c}{1+3c}\kappa\varepsilon_{J}^{\ast}$
$\leq\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}+\tfrac{1+2c}{1+3c}\kappa x.$
Consequently, recalling the definition of $\kappa^{\prime}$ and noting that $2\kappa x\geq\kappa\varepsilon_{J-1}$ since $x\geq\varepsilon_{J}=\varepsilon_{J-1}/2$, we obtain
$\mathbb{P}(\|\bar{f}-f^{\ast\ast}\|_{L_{2}(\mathbb{P}_{X})}>\kappa^{\prime}x)$ $\leq\mathbb{P}(\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}+\tfrac{1+2c}{1+3c}\kappa x>\kappa^{\prime}x)$
$=\mathbb{P}(\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}>2\kappa x)$
$\leq\mathbb{P}(\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}>\kappa\varepsilon_{J-1})$
$=\mathbb{P}(B_{J-1})$
$\leq 6\exp(-n\varepsilon_{J-1}^{2}/2)$
$\leq 6\exp(-nx^{2}/2).$
The last few lines used the definition of the event $B_{J-1}$ and then that $x<\varepsilon_{J-1}.$
Part 5: Bounding $\mathbb{E}_{\vec{X},\vec{Y}}\|\bar{f}-f^{\ast\ast}(\vec{X},\vec{Y})\|_{L_{2}(\mathbb{P}_{X})}^{2}.$
We integrate the bound from Part 4:
$\mathbb{E}_{\vec{X},\vec{Y}}\|\bar{f}-f^{\ast\ast}(\vec{X},\vec{Y})\|_{L_{2}(\mathbb{P}_{X})}^{2}$
$=\int_{0}^{\infty}\mathbb{P}(\|\bar{f}-f^{\ast\ast}\|_{L_{2}(\mathbb{P}_{X})}^{2}>x)\mathrm{d}x$
$=2{\kappa^{\prime}}^{2}\int_{0}^{\infty}t\cdot\mathbb{P}(\|\bar{f}-f^{\ast\ast}\|_{L_{2}(\mathbb{P}_{X})}>\kappa^{\prime}t)\mathrm{d}t$
$\leq 2{\kappa^{\prime}}^{2}\int_{0}^{\varepsilon_{J}^{\ast}}t\mathrm{d}t+2{\kappa^{\prime}}^{2}\int_{\varepsilon_{J^{\ast}}}^{\infty}t\cdot\mathbb{P}(\|\bar{f}-f^{\ast\ast}\|_{L_{2}(\mathbb{P}_{X})}>\kappa^{\prime}t)\mathrm{d}t$
$\leq{\kappa^{\prime}}^{2}\varepsilon_{J^{\ast}}^{2}+12{\kappa^{\prime}}^{2}\int_{\varepsilon_{J^{\ast}}}^{\infty}t\cdot\exp(-nt^{2}/2)\mathrm{d}t$
$={\kappa^{\prime}}^{2}\varepsilon_{J^{\ast}}^{2}+12{\kappa^{\prime}}^{2}\cdot n^{-1}\exp(-n\varepsilon_{J^{\ast}}^{2}/2).$
Our assumption (3.4) implies $n^{-1}\leq\varepsilon_{J^{\ast}}^{2}/\log 2$, while $\exp(-n\eta_{J^{\ast}}^{2}/2)\leq 1$, so the entire bound bounded by $\varepsilon_{J^{\ast}}^{2}$ up to constants depending on $c$, $\sigma$, and $L$. ∎
@@
The $\varepsilon^{\ast}=0$ edge case implies $d=0$ and we trivially achieve the minimax rate of 00. Assume henceforth $\varepsilon^{\ast}>0$.
@@
Case 1: Suppose $n{\varepsilon^{\ast}}^{2}>8\log 2$. Set $\alpha=\min(\sigma/2,1)\in(0,1]$ and $\delta^{\ast}=\alpha\varepsilon^{\ast}$. Note that $\alpha^{2}\leq\frac{\sigma^{2}}{4}$. Then using the monotonicity of $\mathcal{M}_{\mathcal{F}}^{\operatorname{loc}}(\cdot,c)$, we obtain
$\log\mathcal{M}_{\mathcal{F}}^{\operatorname{loc}}(\delta^{\ast},c)$ $\geq\lim_{t\downarrow 0}\log\mathcal{M}_{\mathcal{F}}^{\operatorname{loc}}(\varepsilon^{\ast}-t,c)\geq\lim_{t\downarrow 0}n({\varepsilon^{\ast}}-t)^{2}=\frac{n{\varepsilon^{\ast}}^{2}}{2}+\frac{n{\varepsilon^{\ast}}^{2}}{2}$
$>\frac{n{\delta^{\ast}}^{2}}{2\alpha^{2}}+4\log 2\geq 4\cdot\frac{n{\delta^{\ast}}^{2}}{2\sigma^{2}}+4\log 2\geq 4\left(\frac{n{\delta^{\ast}}^{2}}{2\sigma^{2}}\vee\log 2\right).$
@@
Thus, the lower bound condition from Lemma 2.3 is satisfied for $\delta^{\ast}$, so indeed ${\varepsilon^{\ast}}^{2}$ is a lower bound for the minimax rate up to absolute constants.
@@
We now show ${\varepsilon^{\ast}}^{2}$ is an upper bound. Following [Shrotriya and Neykov, 2022, Theorem 12], we set $\eta=\frac{c}{\sqrt{\Lambda(c/4-1,\sigma,L)}}\wedge 1\in(0,1)$, pick $D$ sufficiently large such that $D\eta>1$, and set $\delta=2D\varepsilon^{\ast}$. Again using the definition of $\varepsilon^{\ast}$ as a supremum and monotonicity of $\mathcal{M}_{\mathcal{F}}^{\operatorname{loc}}(\cdot,c)$, we have
$n\delta^{2}$ $=4D^{2}n{\varepsilon^{\ast}}^{2}=\frac{4}{\eta^{2}}\cdot n(D\eta\cdot\varepsilon^{\ast})^{2}>\frac{4}{\eta^{2}}\cdot\log\mathcal{M}_{\mathcal{F}}^{\operatorname{loc}}(D\eta\varepsilon^{\ast},c)$
@@
$\geq 4\log\mathcal{M}_{\mathcal{F}}^{\operatorname{loc}}(D\eta\varepsilon^{\ast},c)\geq 4\log\mathcal{M}_{\mathcal{F}}^{\operatorname{loc}}\left(\frac{cD\varepsilon^{\ast}}{\sqrt{\Lambda(c/4-1,\sigma,L)}},c\right)$
@@
$=4\log\mathcal{M}_{\mathcal{F}}^{\operatorname{loc}}\left(\frac{c\cdot\delta}{2\cdot\sqrt{\Lambda(c/4-1,\sigma,L)}},c\right).$
Further, since $n{\varepsilon^{\ast}}^{2}>8\log 2$ and clearly $D>1$, we have $n\delta^{2}=4nD^{2}{\varepsilon^{\ast}}^{2}>32D^{2}\log 2>\log 2$, so indeed
@@
$n\delta^{2}>2\log\left[\mathcal{M}_{\mathcal{F}}^{\operatorname{loc}}\left(\frac{c\cdot\delta}{2\sqrt{\Lambda(c/4-1,\sigma,L)}},c\right)\right]^{2}\vee\log 2.$
In other words, $\delta$ satisfies the condition in (3.4), noting by Remark 3.14 we may swap out $2c$ with $c$.
Now consider the non-decreasing map
$\psi\colon 0<x\mapsto nx^{2}-2\log\left[\mathcal{M}_{\mathcal{F}}^{\operatorname{loc}}\left(x\cdot\frac{c}{2\sqrt{\Lambda(c/4-1,\sigma,L)}},c\right)\right]^{2}\vee\log 2.$
@@
We have shown that $\psi(\delta)>0$. Recall our definition of $\varepsilon_{J^{\ast}}$ from Theorem 3.13, and let us show that $\delta\geq\varepsilon_{J^{\ast}}/2$. Suppose no $J\geq 1$ exists such that $\varepsilon_{J}$ satisfies (3.4), in which case $J^{\ast}=1$. Then $\psi(\varepsilon_{J^{\ast}}/2)\leq\psi(\varepsilon_{J^{\ast}})<0<\psi(\delta)$ so $\delta\geq\varepsilon_{J^{\ast}}/2$. On the other hand, suppose some $J$ does exist such that $\varepsilon_{J}$ satisfies (3.4). Because $J^{\ast}<\infty$ is chosen maximally in our algorithm, $\varepsilon_{J^{\ast}+1}=\varepsilon_{J^{\ast}}/2$ does not satisfy (3.4), so we have $\psi(\varepsilon_{J^{\ast}}/2)=\psi(\varepsilon_{J^{\ast}+1})<0<\psi(\delta)$, so $\delta\geq\varepsilon_{J^{\ast}}/2$.
@@
Since Theorem 3.13 upper bounds the minimax rate by $\varepsilon_{J^{\ast}}^{2}$ and we just showed $\delta\gtrsim\varepsilon_{J^{\ast}}$, the rate is also upper bounded by $\delta^{2}\asymp{\varepsilon^{\ast}}^{2}$. Thus, the minimax has a lower and upper bound of rate ${\varepsilon^{\ast}}^{2}$ up to absolute constants. But if ${\varepsilon^{\ast}}^{2}$ is a lower bound, then so is ${\varepsilon^{\ast}}^{2}\wedge d^{2}$; further, the squared diameter is always an upper bound for the minimax rate, so ${\varepsilon^{\ast}}^{2}\wedge d^{2}$ is also an upper bound up to absolute constants. Thus, our minimax rate is ${\varepsilon^{\ast}}^{2}\wedge d^{2}$ up to absolute constants.
Case 2: Suppose $n{\varepsilon^{\ast}}^{2}\leq 8\log 2$. Observe by definition of $\varepsilon^{\ast}$ as a supremum that
$\log\mathcal{M}_{\mathcal{F}}^{\operatorname{loc}}(2\varepsilon^{\ast},c)\leq 4n{\varepsilon^{\ast}}^{2}\leq 32\log 2.$ (B.9)
@@
Now repeating the logic of Prasadan and Neykov [2024, Lemma 1.3] for star-shaped sets, there exists a line segment in $\mathcal{F}$ of length $d/3$ in the $L_{2}(\mathbb{P}_{X})$ distance. Form a ball $B$ of radius $2\varepsilon^{\ast}$ centered at the midpoint of this line segment. Consider the diameter of $B$ that coincides with the chosen line segment in $\mathcal{F}$ at infinitely many points. Since an interval of length $4\varepsilon^{\ast}$ if partitioned into subintervals of length less than $(2\varepsilon^{\ast}/c)$ is composed of at least $2c$ subintervals, for sufficiently large $c$, at least $\exp(32\log 2)$ many points at least $(2\varepsilon^{\ast}/c)$ distance apart can be chosen along this particular diameter of $B$. All of these points cannot also belong to $\mathcal{F}$, otherwise we will have formed a $(2\varepsilon^{\ast}/c)$-packing of $B\cap\mathcal{F}$ with $\exp(32\log 2)$ points, contradicting $\log\mathcal{M}_{\mathcal{F}}^{\operatorname{loc}}(2\varepsilon^{\ast},c)\leq 32\log 2$. This means the line segment of $\mathcal{F}$ we picked is strictly contained in the diameter of $B$ we picked, so that $d/3\leq 4\varepsilon^{\ast}\leq 8\sqrt{2n^{-1}\log 2}$, using (B.9).
@@
Now, consider an $\varepsilon$ which is proportional to $d$. In this case, $n\varepsilon^{2}$ is upper bounded by some constant since $d\leq 24\sqrt{2n^{-1}\log 2}$, while $\log\mathcal{M}_{\mathcal{F}}^{\operatorname{loc}}(2\varepsilon,c)$ can be made arbitrarily large by increasing $c$, to ensure the condition from Lemma 2.3 for the lower bound is attained. So $\varepsilon^{2}$ is a lower bound for the minimax-rate up to constants, and thus by proportionality, $d^{2}$ is a lower bound up to absolute constants. Thus ${\varepsilon^{\ast}}^{2}\wedge d^{2}$ is a lower bound for the minimax rate.